OpenStudy (anonymous):

Let f(x)=4lnIsec(x)+tan(x)I find f"(x)

OpenStudy (anonymous):

For ln(sec(x)) you need to do chain rule. remember that \[\frac{d}{dx}ln(f(x))=\frac{1}{f(x)}*f'(x)\] For the derivative for sec(x), you should note that sec(x) = 1/cos(x) whereupon you can do quotient rule \[\frac{d}{dx}\frac{u}{v}=\frac{uv'+u'v}{v^2}\] Similarly with tan(x) = sin(x)/cos(x) you can apply quotient rule here again. The second derivative isn't much more strenuous.